# OBLR-PO: A Theoretical Framework for Stable Reinforcement Learning

### v1

@@ Appendix B Further Analysis
B.1 Optimal Sampling Strategy Under Computational Budget
In this section, we discuss the optimal sampling strategy $\{G_{t}\}_{t=0}^{T-1}$ and $\{N_{t}\}_{t=0}^{T-1}$ under a fixed computational budget. Suppose that sampling a query requires $C_{Q}$ computation and sampling an output requires $C_{O}$ computation. The optimization problem can be formally stated as follows:
$\min_{\{N_{t}\}_{t=0}^{T-1},\{G_{t}\}_{t=0}^{T-1}}\mathbb{E}[\mathscr{L}(\theta_{T})]\quad\text{subject to}\quad\sum_{t=0}^{T-1}\left(C_{Q}N_{t}+C_{O}N_{t}G_{t}\right)\leq C_{1}.$ (102)
Note that $C_{O}\gg C_{Q}$, which allows us to approximate the constraint as
$\sum_{t=0}^{T-1}N_{t}G_{t}\leq C_{2}.$ (103)
Thus, the optimization problem can be rewritten as
$\min_{\{N_{t}\}_{t=0}^{T-1},\{G_{t}\}_{t=0}^{T-1}}\mathbb{E}[\mathscr{L}(\theta_{T})]\quad\text{subject to}\quad\sum_{t=0}^{T-1}N_{t}G_{t}\leq C_{2}.$ (104)
@@ B.1 Optimal Sampling Strategy Under Computational Budget
Under the computational budget constraint (103), the optimal sampling strategy is given by
$N_{t}G_{t}=\frac{C_{2}+\sum_{t=0}^{T-1}\frac{\text{tr}(\mathbf{H}(\theta_{t}))}{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}}{\sum_{t=0}^{T-1}\sqrt{\text{tr}(\mathbf{H}(\theta_{t}))}}\sqrt{\text{tr}(\mathbf{H}(\theta_{t}))}-\frac{\text{tr}(\mathbf{H}(\theta_{t}))}{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}.$ (105)
@@
$\sum_{t=0}^{T-1}\frac{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\text{tr}(\mathbf{H}(\theta_{t}))},$ (106)
@@
$\sum_{t=0}^{T-1}\frac{\frac{1}{N_{t}G_{t}}\text{tr}(\mathbf{H}(\theta_{t}))\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\text{tr}(\mathbf{H}(\theta_{t}))}=\sum_{t=0}^{T-1}\frac{\text{tr}(\mathbf{H}(\theta_{t}))}{N_{t}G_{t}+\frac{\text{tr}(\mathbf{H}(\theta_{t}))}{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}}.$ (107)
@@
$\left(\sum_{t=0}^{T-1}\left(N_{t}G_{t}+\frac{\text{tr}(\mathbf{H}(\theta_{t}))}{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}\right)\right)\cdot\left(\sum_{t=0}^{T-1}\frac{\text{tr}(\mathbf{H}(\theta_{t}))}{N_{t}G_{t}+\frac{\text{tr}(\mathbf{H}(\theta_{t}))}{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}}\right)\geq\left(\sum_{t=0}^{T-1}\sqrt{\text{tr}(\mathbf{H}(\theta_{t}))}\right)^{2}.$ (108)
@@
$\frac{\sqrt{\text{tr}(\mathbf{H}(\theta_{t}))}}{N_{t}G_{t}+\frac{\text{tr}(\mathbf{H}(\theta_{t}))}{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}}=\operatorname{Const}.$ (109)
Substituting this into the computational budget constraint (103) yields the result. ∎

### v2

@@ Appendix B Further Analysis
B.1 Optimal Query Sampling Strategy Under a Data Constraint
In this section, we study the optimal query sampling schedule $\{N_{t}\}_{t=0}^{T-1}$ under a fixed data budget. Formally, we consider
$\min_{\{N_{t}\}_{t=0}^{T-1},\,\{G_{t}\}_{t=0}^{T-1}}\;\mathbb{E}\big[\mathscr{L}(\theta_{T})\big]\quad\text{s.t.}\quad\sum_{t=0}^{T-1}N_{t}\leq C.$ (110)
@@ B.1 Optimal Query Sampling Strategy Under a Data Constraint
Under the computational budget constraint, the optimal sampling strategy is given by
$N_{t}=\frac{C+\sum_{t=0}^{T-1}\frac{\text{tr}(\boldsymbol{H}(\theta_{t}))}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}}{\sum_{t=0}^{T-1}\sqrt{\text{tr}(\boldsymbol{H}(\theta_{t}))}}\sqrt{\text{tr}(\boldsymbol{H}(\theta_{t}))}-\frac{\text{tr}(\boldsymbol{H}(\theta_{t}))}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}.$ (111)
@@
$\sum_{t=0}^{T-1}\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\text{tr}(\boldsymbol{H}(\theta_{t}))},$ (112)
@@
$\sum_{t=0}^{T-1}\frac{\frac{1}{N_{t}}\text{tr}(\boldsymbol{H}(\theta_{t}))\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\text{tr}(\boldsymbol{H}(\theta_{t}))}=\sum_{t=0}^{T-1}\frac{\text{tr}(\boldsymbol{H}(\theta_{t}))}{N_{t}+\frac{\text{tr}(\boldsymbol{H}(\theta_{t}))}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}}.$ (113)
@@
$\left(\sum_{t=0}^{T-1}\left(N_{t}+\frac{\text{tr}(\boldsymbol{H}(\theta_{t}))}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}\right)\right)\cdot\left(\sum_{t=0}^{T-1}\frac{\text{tr}(\boldsymbol{H}(\theta_{t}))}{N_{t}+\frac{\text{tr}(\boldsymbol{H}(\theta_{t}))}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}}\right)\geq\left(\sum_{t=0}^{T-1}\sqrt{\text{tr}(\boldsymbol{H}(\theta_{t}))}\right)^{2}.$ (114)
@@
$\frac{\sqrt{\text{tr}(\boldsymbol{H}(\theta_{t}))}}{N_{t}+\frac{\text{tr}(\boldsymbol{H}(\theta_{t}))}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}}=\operatorname{Const}.$ (115)
Substituting this into the computational budget constraint yields the result. ∎